# Probabilities of Chat LLMs Are Miscalibrated but Still Predict Correctness on Multiple-Choice Q&A

### v3

@@ 7 Conclusion
In this paper, we showed that the MSPs of chat LLMs are miscalibrated but still provide a reliable signal of uncertainty. Furthermore, the reliability of this signal improves as model capabilities improve, but the same is not true of calibration. Several specific directions for future work are discussed in Section 8. Overall, we are excited about developing more robust methods for mistake detection in LLMs for a variety of contexts.
8 Limitations
@@
Our study has several important limitations. One is the restriction to multiple-choice questions, which simplifies the problem in several ways. First, it removes the need to distinguish between multiple correct answers (“aleatoric uncertainty”) and no good answers (“epistemic uncertainty”), both of which could result in low MSPs. The restriction to multiple-choice also enabled us to tie the LLM’s answer to a single token and thus a single MSP. Future work could handle free-response questions by aggregating MSPs across tokens in a clever way. A further challenge could be multi-step decision-making problems which may involve aggregating uncertainty not only across multiple tokens in a single response, but across multiple responses on different time steps.
@@
Another limitation is our reliance on labeled data to transform our scientific insights into a practical method for abstention. We only used 20 data points, and labeled data was only used to choose the threshold, but a fully unsupervised method would be advantageous in many settings. However, we remind the reader that our primary contribution is the scientific finding of good correctness prediction despite miscalibration, not the proof-of-concept abstention experiments in Section 6.
Third, we only studied fine-tuned “chat” models and did not include non-fine-tuned base models in our experiments. As mentioned in Section 3, this choice was due to the failure of the base models to consistently respond in the correct format. Nevertheless, the exclusion of the base models means that our experiments do not determine the impact of the fine-tuning process: we only characterize properties of the resulting fine-tuned models. As discussed in Section 2, Kadavath et al. (2022) do study the calibration of base models, but they only study a single family of models. Our guess is that that particular family of base models is more reliable in terms of response format.
Lastly, our study is entirely within the artificial setting of multiple choice Q&A benchmarks. Readers should not assume that our method will work equally well on realistic user queries (and even in the artificial setting, the methods we study are not perfect). We would not want a reader implementing method to assume that they no longer need to worry about false responses; this is one potential risk from our work.
9 Ethical considerations
Misleading answers from LLM can have significant harmful consequences. We hope that our work constitutes a step in direction of averting such responses.
We do not anticipate any direct ethical consequences of our work. Perhaps the most plausible way our work could have negative consequences is if a reader mistakenly thinks that using our method will completely eliminate false LLM responses and relies too heavily on LLM responses as a result. We believe that we have appropriately disclosed and discussed that risk in Section 8.
@@ Appendix A Details on main experiments
Here we include details on the methodology and results of our main experiments. First, we include the second prompt phrasing we used (Figure 5). Recall that Figure 1 shows the first phrasing.
You will be asked a multiple-choice question. Respond with the letter which corresponds to the correct answer, followed by a period. There is no need to provide an explanation, so your response should be very short. Now here is the question: In the nitrogen cycle, nitrogen can return to the lithosphere directly from the atmosphere by
A. lightning.
B. cellular respiration.
C. air pollution.
D. condensation.
Answer:
Figure 5: The second prompt phrasing we used.
@@ A.1 Details on calibration results
LLM Q&A accuracy Calibration error (×$\times 100$)
Falcon 7B 29.4 7.8
Falcon 40B 43.8 15.1
Llama 2 7B 40.4 32.2
Llama 2 70B 58.1 32.2
Llama 3.0 8B 58.5 22.5
Llama 3.0 70B 78.4 22.4
Llama 3.1 8B 62.8 12.5
Llama 3.1 70B 80.8 9.2
Mistral 7B 56.3 38.4
Mixtral 8x7B 69.0 29.8
SOLAR 10.7B 67.1 25.5
Yi 6B 53.0 19.5
Yi 34B 70.1 11.5
GPT-3.5 Turbo 67.5 22.1
GPT-4o 86.6 16.7
Table 3: Q&A accuracy and calibration error (scaled by 100 for readability) for each model.
@@
As mentioned in Section 4, the each model’s calibration error was computed by first computing the calibration error for each combination of model-dataset-phrasing, and then averaging over datasets and phrasings to obtain a per-model total error. This was done to avoid downweighting datasets with fewer questions. This data is shown in Figure 3 in Section 4 and also in Table 3 here in this appendix.
@@
However, this approach does not make sense for the calibration curves in Figure 2. This is because calibration curves are not averages: they are obtained by bucketing each data point (in our case, a question-response pair is a data point). In order to avoid downweighting the smaller datasets, we duplicated data points from those datasets so that the total number of points per dataset was roughly 6000. Concretely, we duplicated data points from TruthfulQA by $6000/817\approx 7$ and from ARC by $6000/2590\approx 2$. Although this solution has drawbacks, we felt that it was the least bad option. The data in Figure 2 uses this duplication approach.
For the curious reader, Figure 6 provides version where the data is not duplicated (and as a result, downweights TruthfulQA and ARC). Qualitatively, there are no significant differences between this figure and the duplication approach in Figure 2.
Figure 6: A variant of Figure 2 where smaller datasets are not duplicated and thus are downweighted.
Q&A accuracy MSP AUROC Max Logit AUROC
ARC-Challenge 69.5 71.9 67.2
HellaSwag 58.3 67.8 62.2
MMLU 54.4 68.5 64.1
TruthfulQA 45.8 66.6 62.0
WinoGrande 59.3 57.9 54.5
Table 4: Q&A accuracy and AUROCs per dataset. All values are percentages, averaged over all models and prompt phrasings.
A.2 Details on AUROC results
Statistical test details.
We discussed the basic statistical test setup in Section 5. Here we make two additional comments:
  1. 1.
    As the non-parametric equivalent of the $t$-test, MWU requires the data to be independent but not normally distributed, which our randomly-sampled multiple choice questions satisfy.
  2. 2.
    We chose the significance threshold of $\alpha=10^{-4}$ to account for a Bonferroni correction (Bonferroni, 1936), which is applied when performing multiple hypothesis tests (in our case, 280) to ensure that the chance of falsely rejecting any null hypothesis is small. Starting from the standard threshold of $\alpha=0.05$, the Bonferroni correction yields $\alpha=0.05/280\approx 1.8\times 10^{-4}$. We adhere to the stricter threshold of $1\times 10^{-4}$ for simplicity. Since there are far fewer cross-model comparisons (e.g., Figures 3 and 4).
WinoGrande.
WinoGrande was by far the hardest dataset for our correctness prediction task (Table 4). Our best hypothesis for this discrepancy is that WinoGrande is intentionally adversarial and tries to “trick" the model. An illustrative question from this dataset is “Neil told Craig that he has to take care of the child for the day because     did it last time." Even for some humans, it could be unclear whether Neil is assuming responsibility or assigning responsibility. One wrinkle is that the Q&A accuracy on WinoGrande is comparable to other datasets, so it is not the case that this dataset is “harder” in general: it is harder only for predicting correctness. Regardless, despite WinoGrande’s average MSP AUROC of 57.9%, Llama 3.1 70B and GPT-4o Turbo still achieved AUROCs of 75.8% and 78.3% respectively on this dataset (Table 12), suggesting that this difficulty is surmountable for sufficiently capable models.
Minimal correlation between model size and AUROC.
The PCCs between model size and AUROC were $r=0.57$ ($p=0.04$) and $r=0.36$ ($p=0.23$) for MSP and Max Logit, respectively (Figure 7). It is unsurprising that some correlation exists, due to the known correlation between size and model capabilities (e.g., Q&A accuracy) and our correlation between Q&A accuracy and AUROC. However, the relatively weak correlation between model size and AUROC may suggest that adding more parameters does not directly improve predictive power of these classifiers outside of improving the model’s overall capabilities. The OpenAI models were excluded from these charts because we do not know their size.
Figure 7: AUROC vs model size for MSP (left) and Max Logit (right). The PCCs for MSP and Max Logit were $r=0.57$ ($p=0.04$) and $r=0.36$ ($p=0.23$) respectively. GPT-3.5 Turbo and GPT-4o were excluded since their sizes are unknown.
Figure 8: Average AUROC vs Q&A accuracy based on prompt phrasing (the two phrasings can be found in Figures 1 and 5). All values are averaged over the five datasets.
Impact of prompt phrasing.
The two different prompt phrasings (Figures 1 and 5) did not yield significantly different results (Figure 8). This suggests that our results are robust to minor modifications to prompt phrasing.
A.3 Details on Q&A-with-abstention results
@@
Figure 9 is based on the same data as Table 2, but shows each model’s performance across the entire range of possible thresholds. A threshold of zero corresponds to the base LLM and the black dot indicates the threshold chosen during the training phase (using 20 labeled data points), which is also the threshold used to compute the score in Table 2 using 20 data points. One can see that the chosen thresholds are not quite optimal, but 20 data points was still enough to produce substantial improvements over the baseline of not abstaining.
@@
Table 5 presents the average abstention frequencies across all datasets, corresponding to the scores from Table 2. We also provide dataset-level versions of Table 5 in Appendix F.
@@
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B 00 $89.9$ $98.4$ 00 $89.9$ $98.4$
Falcon 40B 00 $90.4$ $98.4$ 00 $94.4$ $98.4$
Llama 2 7B 00 $21.6$ $32.7$ 00 $98.0$ $99.1$
Llama 2 70B 00 $15.4$ $11.9$ 00 $36.4$ $64.2$
Llama 3.0 8B 00 $31.2$ $58.1$ 00 $52.0$ $58.1$
Llama 3.0 70B 00 $0.0$ $0.0$ 00 $42.4$ $26.5$
Llama 3.1 8B 00 $40.8$ $33.9$ 00 $58.2$ $65.6$
Llama 3.1 70B 00 $15.6$ $0.0$ 00 $15.6$ $17.7$
Mistral 7B 00 $38.1$ $8.2$ 00 $41.8$ $74.2$
Mixtral 8x7B 00 $9.3$ $27.1$ 00 $9.3$ $27.1$
SOLAR 10.7B 00 $0.0$ $0.0$ 00 $5.8$ $15.2$
Yi 6B 00 $52.7$ $44.1$ 00 $77.0$ $98.1$
Yi 34B 00 $5.0$ $3.3$ 00 $5.0$ $3.3$
GPT-3.5 Turbo 00 $15.3$ 00 $45.6$
GPT-4o 00 $2.2$ 00 $2.2$
Table 5: Frequency of abstention in the Section 6 experiments.
Finally, Tables 6 and 7 are analogues of Table 2 using 10 data points and half of all data points for training, respectively.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B $-41.2$ $-0.4$ $-1.0$ $-111.8$ $-5.7$ $-1.6$
Falcon 40B $-12.4$ $-7.4$ $-8.2$ $-68.6$ $-14.7$ $-16.8$
Llama 2 7B $-19.3$ $-11.0$ $-17.8$ $-79.0$ $-3.0$ $-8.7$
Llama 2 70B $16.2$ $21.3$ $19.1$ $-25.7$ $7.4$ $3.0$
Llama 3.0 8B $17.1$ $22.6$ $17.2$ $-24.4$ $10.4$ $8.3$
Llama 3.0 70B $56.8$ $56.8$ $56.8$ $35.2$ $47.3$ $35.2$
Llama 3.1 8B $25.7$ $25.8$ $26.0$ $-11.5$ $3.8$ $-8.8$
Llama 3.1 70B $61.7$ $62.9$ $61.7$ $42.5$ $51.5$ $42.5$
Mistral 7B $12.6$ $16.9$ $15.5$ $-31.0$ $-6.8$ $-4.2$
Mixtral 8x7B $38.1$ $38.8$ $35.4$ $7.2$ $12.6$ $14.9$
SOLAR 10.7B $34.2$ $34.2$ $34.2$ $1.3$ $1.3$ $9.9$
Yi 6B $6.0$ $15.5$ $10.4$ $-41.0$ $6.5$ $1.1$
Yi 34B $40.2$ $40.3$ $40.3$ $10.2$ $14.3$ $10.8$
GPT-3.5 Turbo $35.0$ $38.5$ $2.5$ $27.3$
GPT-4o $73.2$ $73.3$ $59.9$ $65.8$
Table 6: Q&A with abstention results for $k=10$. See Table 2 for an explanation of the scoring scheme.
Balanced Conservative
LLM Base MSP Max Logit Base MSP Max Logit
Falcon 7B $-41.0$ $0.0$ $0.0$ $-111.5$ $-0.1$ $0.0$
Falcon 40B $-12.4$ $1.9$ $0.2$ $-68.5$ $0.0$ $0.0$
Llama 2 7B $-19.1$ $0.0$ $0.3$ $-78.7$ $0.0$ $-0.1$
Llama 2 70B $16.9$ $23.5$ $20.2$ $-24.7$ $7.3$ $4.8$
Llama 3.0 8B $17.1$ $23.9$ $23.3$ $-24.4$ $12.6$ $11.2$
Llama 3.0 70B $56.1$ $57.8$ $56.5$ $34.2$ $48.4$ $41.9$
Llama 3.1 8B $25.5$ $30.3$ $26.9$ $-11.7$ $17.7$ $13.1$
Llama 3.1 70B $61.4$ $62.8$ $61.4$ $42.1$ $53.8$ $42.6$
Mistral 7B $11.4$ $16.2$ $15.3$ $-32.8$ $0.0$ $4.9$
Mixtral 8x7B $37.4$ $38.4$ $37.8$ $6.0$ $13.5$ $13.6$
SOLAR 10.7B $33.7$ $34.5$ $34.5$ $0.6$ $8.4$ $13.5$
Yi 6B $5.0$ $14.6$ $9.7$ $-42.4$ $6.1$ $1.1$
Yi 34B $40.4$ $41.6$ $40.4$ $10.6$ $18.5$ $19.1$
GPT-3.5 Turbo $34.8$ $38.9$ $2.1$ $28.4$
GPT-4o $73.4$ $73.9$ $60.1$ $66.1$
Table 7: Q&A with abstention results for $k=$ half of all data points. See Table 2 for an explanation of the scoring scheme.
Figure 10: Calibration results for 5-shot prompting. Left: calibration curves for the 15 models. Right: as for zero-shot, there is no correlation between calibration error and average Q&A accuracy $(r=-0.02,p=0.95)$.
Figure 11: AUROC results for 5-shot prompting. MSP AUROC (left) still exhibits a strong correlation with Q&A accuracy $(r=0.62,p<0.005)$. However, Max Logit AUROC (right) no longer exhibits a statistically significant correlation with average Q&A accuracy $(r=0.43,p=0.14)$.
Appendix B Results for 5-shot prompting
As discussed in Section 3, we chose zero-shot prompting for our main results in order to test our hypothesis in the simplest setting possible. In this section, we show how the results change if 5-shot prompting is used instead. The short answer is that the same correlations and trends hold, except for Max Logit. The AUROC values are also slightly lower overall. This could suggest that few-shot prompting partially disrupts the innate uncertainty information in LLMs, but we do not have the evidence to conclude this definitively.
Figure 10 presents the calibration results for 5-shot prompting. As in the zero-shot setting, models are consistently overconfident and there is no correlation between calibration error and Q&A accuracy $(r=-0.02,p=0.95)$. If anything, calibration error is slightly higher in the 5-shot setting, but this may just be noise.
Figure 11 presents the AUROC correctness prediction results for 5-shot prompting. The correlation between MSP AUROC and Q&A accuracy remains $(r=0.69,p<0.005)$, but Max Logit AUROC no longer exhibits a statistically significant correlation with Q&A accuracy $(r=0.43,p=0.14)$.
Appendix C What if we use the product of MSPs across tokens?
As discussed in Section 3, an LLM response typically consists of multiple tokens, each with its own MSP. For our main results, we selected the MSP of the token which indicates the LLM’s answer, i.e., “A”/“B”/etc. We chose this process because we believe that it is the fairest way of computing the probability that the model assigns to its answer. However, it is important to know whether our results would change if the MSP were computed differently. In this section, we consider one potential alternative: computing the product of MSPs across all tokens. This value corresponds to the overall probability the model assigns to its output. [^14] Since the product of maximum logits does not admit such an interpretation, we only study the product of MSPs here. Does this method produce different results?
The answer is generally no, for both calibration (Figure 12) and correctness prediction (Figure 13). The biggest difference is that there the product of MSPs is typically smaller than the MSP of the answer token, which is expected. As a result, the LLMs are no longer consistently overconfident. However, the models remain consistently miscalibrated: they are now simply underconfident in addition to being overconfident. (The underconfident instances may correspond to responses where the LLM included an explanation of its answer, which would cause the product of MSPs to be artificially low.) Also, there remains no correlation between calibration error and Q&A accuracy $(r=-0.19,p=0.50)$.
With regards to correctness prediction, we find the same correlation between AUROC and Q&A accuracy with approximately the same strength $(r=0.83,p=0.00014)$.
These results show that our findings are robust to modifications in how the MSP is computed across tokens.
Figure 12: Calibration results for the product of MSPs. There is still no evidence for a correlation between calibration error and Q&A accuracy $(r=-0.19,p=0.50)$.
Figure 13: AUROC results for the product of MSPs. There remains a strong correlation between AUROC and Q&A accuracy $(r=0.83,p<10^{-3})$.
Appendix D Details on appropriate use of models and datasets
D.1 Licenses
The Hugging Face transformers library uses the Apache 2.0 license. The licenses for the datasets we used are:
  1. 1.
    ARC-Challenge: CC BY-SA 4.0
  2. 2.
    HellaSwag: MIT
  3. 3.
    MMLU: MIT
  4. 4.
    TruthfulQA: Apache 2.0
  5. 5.
    WinoGrande: CC-BY
The licenses for the models we used are:
  1. 1.
    Falcon 7B and 40B: Apache 2.0
  2. 2.
    Llama 2 7B and 70B: Llama 2 Community License Agreement
  3. 3.
    Llama 3 8B and 70B: Llama 3 Community License Agreement
  4. 4.
    Llama 3.1 8B and 70B: Llama 3.1 Community License Agreement
  5. 5.
    Mistral 7B and Mixtral 8x7B: Apache 2.0
  6. 6.
    SOLAR 10.7B: CC BY-NC 4.0
  7. 7.
    Yi 6B and 34B: Yi Series Models Community License Agreement
  8. 8.
    GPT-3.5 Turbo and GPT-4o: OpenAI Terms of Service
All models and datasets were used in a manner consistent with their licenses and intended use.
D.2 Dataset contents
As mentioned, we used the established datasets from the original Hugging Face LLM leaderboard (Beeching et al., 2023), so we did not have control over their contents. These datasets have each been curated to some extent, and we did not find any instances of personally identifiable information (PII) or toxic content. However, we are not aware of any exhaustive efforts to remove PII or toxic content. Ideally, we would use datasets which are guaranteed to not include PII or toxic content, but we felt that using the established datasets from said leaderboard was more important.
All datasets are in English. They cover a variety of domains such as math, science, history, and common sense reasoning. We are not aware of efforts to document precisely which demographic groups are covered.
Usage of AI in our research.
AI assistants were used in three ways in our work: (1) basic writing support such as improving phrasing, (2) feedback on the overall paper, (3) coding assistance.
Appendix E Caveat for Falcon 7B
Initially, many of Falcon 7B’s responses fell into the “unparseable" category described in Section 3. Upon investigation, we found that many of these responses were simply a period or an end-of-text token. Removing the final newline in the prompt resolved this behavior, so we believe that this newline was somehow convincing the model that was conversation was “over". These initial results had the side effect of making it very easy to detect wrong answers, since a solitary period is obviously not a correct answer. For this reason, we removed the final newline for Falcon 7B only. We considered removing the final newline for all models or excluding Falcon 7B entirely, but we felt that our chosen approach would be more scientifically honest. As Falcon 7B performed by far the worst on both Q&A and AUROC even with this concession, we do not think this decision holds much import, but we report it for transparency.
Appendix F Dataset-level results
For the curious reader, here we present dataset-level versions of Table 1 (AUROC and Q&A accuracy), Table 2 (results for Q&A with abstention), and Table 5 (frequency of abstention on Q&A with abstention experiments).

### v4

@@ 8 Conclusion
In this paper, we provided a thorough evaluation of correctness awareness in LLMs. We showed that the MSPs of chat LLMs are miscalibrated but still provide a reliable signal of uncertainty. Furthermore, for chat LLMs, the reliability of this signal improves as model capabilities improve, but the same is not true of calibration. This contrasts with base LLMs, where correctness prediction and calibration improve in tandem. Our results pinpoint chat fine-tuning as the source of this divergence.
@@
Our study has several limitations. One is the restriction to multiple-choice questions, which simplifies the problem in several ways. First, it removes the need to distinguish between multiple correct answers (“aleatoric uncertainty”) and no good answers (“epistemic uncertainty”), both of which could result in low MSPs. The restriction to multiple-choice also enabled us to tie the LLM’s answer to a single token and thus a single MSP. Future work could handle free-response questions by aggregating MSPs across tokens in a clever way. A further challenge could be multi-step decision-making problems which may involve aggregating uncertainty not only across multiple tokens in a single response, but also across multiple responses on different time steps.
@@
Another limitation is our reliance on labeled data to transform our scientific insights into a practical method for abstention. We only used 20 data points, and labeled data was only used to choose the threshold, but a fully unsupervised method would be advantageous in many settings.
We would also like to better understand when and why these methods fail. Are there particular subcategories of unfamiliar situations that are especially challenging to identify? For example, why was the WinoGrande dataset so much harder for our correctness prediction task?
More broadly, we are excited about developing more robust methods for mistake detection in LLMs, both for Q&A tasks and other contexts.
Broader impact statement
The capabilities of AI systems have advanced rapidly over the past several years and will likely continue to grow. In order to ensure that AI is beneficial for society, we believe it is paramount to understand and mitigate the risks of such systems. In this paper, we focus on one particular risk: harmful responses from LLMs. We hope that our work contributes to the ongoing efforts to mitigate harmful LLM responses.
We do not think it is likely for our work to inadvertently cause harm, but one possibility is worth mentioning. If a reader were to assume that abstention methods can completely eliminate false responses, that reader might be more likely to fall prey to false responses when they do inevitably still occur. We caution readers to remain vigilant about false responses from LLMs.
@@ Acknowledgements
Author contributions
B. Plaut conceived the project with feedback from N. Khanh and T. Trinh. B. Plaut led the project. B. Plaut designed the initial correctness prediction experiments. N. Khanh suggested adding calibration analysis. All authors contributed to experiment design. B. Plaut implemented and ran all experiments. T. Trinh implemented and ran the statistical analysis. B. Plaut implemented and ran the overall analysis and visualization pipeline. B. Plaut led the writing, with significant contributions from N. Khanh and T. Trinh.
@@ Appendix A Details on main experiments
Here we include details on the methodology and results of our main experiments. First, we include the second prompt phrasing we used (Figure˜8). Recall that Figure˜1 shows the first phrasing.
@@ A.1 Details on calibration results
Table 6: Calibration error and Q&A accuracy for each model.
LLM Q&A accuracy Calibration error ($\times$100)
Falcon 7B 30.1 14.8
Falcon 40B 40.5 11.9
Llama 2 7B 39.7 40.7
Llama 2 70B 58.7 32.8
Llama 3.0 8B 59.1 20.4
Llama 3.0 70B 78.8 16.5
Llama 3.1 8B 62.0 10.8
Llama 3.1 70B 80.6 9.2
Mistral 7B 57.0 40.9
Mixtral 8x7B 69.9 29.6
SOLAR 10.7B 67.5 30.8
Yi 6B 49.6 27.3
Yi 34B 69.1 17.4
GPT-3.5 Turbo 67.5 21.5
GPT-4o 86.9 11.5
@@
As discussed in Section˜4, each model’s calibration error was computed by first computing the calibration error for each combination of model-dataset-phrasing, and then averaging over datasets and phrasings to obtain a per-model total error. This was done to avoid downweighting datasets with fewer questions. This data is shown in Figure˜2 (right) in Section˜4 and also in Table˜6 here in this appendix.
@@
However, this approach does not make sense for the calibration curves in Figure˜2. This is because calibration curves are not averages: they are obtained by bucketing each data point (in our case, a question-response pair is a data point). In order to avoid downweighting the smaller datasets, we duplicated data points from those datasets so that the total number of points per dataset was 6,000. Concretely, we duplicated each data point from TruthfulQA by $7$ to obtain $7\cdot 817=5719$ questions and then randomly sampled an additional 6,0005,7196,000-5,719 questions to obtain 6,000 total questions. Similarly, we duplicated each data point from ARC by $2$, and then randomly sampled $6,000-2\cdot 2,590$ extra questions. Although this solution has drawbacks, we felt that it was the most reasonable option.
A.2 Details on Q&A-with-abstention results
@@
Figure˜9 is based on the same experimental data as Table˜4, but shows each model’s performance across the entire range of possible thresholds. A threshold of zero corresponds to the base LLM and the black dot indicates the threshold chosen during the training phase (using 20 labeled data points per dataset), which is also the threshold used to compute the score in Table˜4. One can see that the chosen thresholds are not quite optimal, but 20 data points was still enough to produce substantial improvements over the baseline of not abstaining.
@@
We can also visualize this by plotting the performance of each model as a function of the amount of training data, as shown in Figure˜10.
Figure 10: Q&A-with-abstention scores as a function of the amount of training data. The x-axis is the number of data points per dataset included in the training data (referred to as $k$ in Section˜6.1). Overall, scores significantly improve as $k$ increases to about 20, but there is minimal improvement if $k$ increases beyond that.
Table 7 shows the average abstention frequencies across all datasets, corresponding to the scores from Table 4.
Table 7: Frequency of abstention in the Section 6 experiments, as percentages.
@@
LLM No abstain MSP Max Logit No abstain MSP Max Logit
Falcon 7B 0 $86.5$ $90.4$ 0 $86.5$ $90.4$
Falcon 40B 0 $93.7$ $68.1$ 0 $94.7$ $99.7$
Llama 2 7B 0 $66.9$ $88.1$ 0 $94.4$ $88.1$
Llama 2 70B 0 $6.7$ $14.2$ 0 $50.1$ $52.2$
Llama 3.0 8B 0 $36.8$ $36.9$ 0 $54.7$ $47.6$
Llama 3.0 70B 0 $10.4$ $0.0$ 0 $17.1$ $19.4$
Llama 3.1 8B 0 $22.9$ $58.9$ 0 $67.2$ $58.9$
Llama 3.1 70B 0 $11.4$ $0.0$ 0 $11.4$ $13.8$
Mistral 7B 0 $4.5$ $29.6$ 0 $100.0$ $92.7$
Mixtral 8x7B 0 $1.4$ $6.2$ 0 $6.7$ $71.1$
SOLAR 10.7B 0 $6.8$ $0.0$ 0 $11.5$ $9.7$
Yi 6B 0 $34.3$ $36.6$ 0 $86.8$ $83.0$
Yi 34B 0 $20.4$ $22.9$ 0 $21.1$ $30.3$
GPT-3.5 Turbo 0 $47.6$ 0 $47.6$
GPT-4o 0 $0.4$ 0 $0.4$
A.3 Details on base models
Table˜8 provides the precise AUROC and Q&A accuracy values in Figure˜7 (analogous to Figure˜3). Table˜9 provides Q&A-with-abstention results for the base models (analogous to Table˜4). As with the chat models, we see that $k=20$ labeled data points per dataset is sufficient to improve over the baseline of never abstaining.
Table 8: AUROC results for the base models. See Table 2 for an explanation of the $p$-values.